# Can LLMs Grasp Implicit Cultural Values? Benchmarking LLMs’ Metacognitive Cultural Intelligence with CQ-Bench

### v1

@@ 3.2 Cultural Intelligence across Different LLMs
Although the attitude detection task should be easier than value selection, its scores are even lower. One reason is that models struggle to distinguish neutral stances, such as neither agree nor disagree. While they can easily differentiate between agree and disagree, the presence of a neutral option can cause confusion. Even when a model correctly identifies agree, it may be distracted by the neutral choice and incorrectly select neither agree nor disagree. Additionally, models find it challenging to differentiate between varying levels of severity, such as not often and not at all often. Although it is also challenging for humans, humans can do better in identifying nuanced attitudes than models. We present results after merging options of varying levels of severity in Table 1. We show option candidates in Appendix A.2.
Smaller models fail on long CoT reasoning in cultural intelligence task.
Although CoT reasoning significantly enhances performance in large models, smaller models often struggle with long CoT reasoning. LLaMA models, in particular, perform poorly in adhering to CoT reasoning across both AD and VS tasks. While Qwen 7B follows CoT reasoning well in AD, its performance declines significantly in VS as the reasoning steps become longer. A manual inspection reveals that its final outputs often consist of random values or irrelevant phrases that fail to focus on the given options.
The DS-distill models exhibit slightly better instruction-following capabilities, outperforming Qwen 7B and LLaMA 8B in VS. While they do not strictly adhere to the prescribed format, their reasoning process generally aligns with the ideas provided in the prompt. However, their CQ reasoning ability remains weaker than their mathematical reasoning skills, resulting in final scores that are still lower than in the no-reasoning setting.
@@
Value extraction requires strong reasoning and summarization capabilities, and we find that smaller models often struggle with this task, frequently producing nonsensical outputs. As a result, we focus our evaluation on four models: Qwen 2.5 32B, GPT-4o-mini, o3-mini, and DeepSeek-V3. We also evaluate a smaller subset of 25 stories using o1 and DeepSeek-R1. Interestingly, unlike attitude detection and value selection—where larger models consistently outperform smaller ones—we observe that a weaker model (Qwen 2.5 32B) can outperform stronger ones like o3-mini and o1. One possible explanation is that current CoT reasoning methods are not well-suited for open-ended generation; models tend to perform better when given predefined options. Nonetheless, DeepSeek-V3 and R1 outperform other models in VE.
3.3 Cultural Intelligence across Different Categories of Culture Values
@@
We present results on category-specific datasets in Figure 3, showing the performance of six models in the no-reasoning setting. Complete results across all models and settings are provided in the Appendix B.1. Overall, the results show that models perform better in understanding political, social, and ethical values, achieving performance that is better than or comparable to the random setting. However, they perform worse in the religious values domain across both tasks. Specifically, while models show stronger performance in attitude detection on the political dataset, they perform better in value selection on the ethical dataset. This suggests that it is easier for models to infer people’s political stances but more challenging for them to identify the specific topics being discussed in political contexts. In contrast, in the ethics domain, models find it easier to identify the main topics being discussed. For value extraction, overall, models perform better on category-specific results except for deepseek-V3, as they only need to cover a single topic, making value extraction easier. Similarly, models performed worse in extracting religious values.
3.4 Improving Cultural Intelligence on Smaller Models
In Section 3.2, we mention that smaller models struggle with long CoT reasoning. Since these models do not follow instructions well, we initially experiment with few-shot prompting to assess whether it can enhance their reasoning ability. However, because each demonstration is lengthy, we use one-shot prompting instead. The results, shown in Figure 4, indicate that while one-shot prompting improves instruction following and enhances performance on some datasets, it does not necessarily improve reasoning ability. In religious dataset AD task, the performance of Qwen 7B one-shot reasoning even decreases.
CQ ability distillation for small models
To further address this issue, we apply supervised fine-tuning on smaller models. Specifically, we distill the reasoning process from o3-mini into smaller models using the random setting dataset consisting of only 500 training samples. After training with LoRA (Hu et al., 2021) for five epochs, we evaluate the models on a category-specific dataset. Since the category-specific dataset contains values not present in the random dataset, it can be considered an out-of-domain dataset approximately. This allows us to assess whether the model has truly learned the reasoning process rather than merely memorizing the values themselves.
We fine-tune four models: Qwen-7B, Qwen-14B, LLaMA 3.1-8B, and LLaMA 3.2-3B, and present the full results alongside o3-mini in Figure 4. Except for the ethical dataset, where SFT results show a slight drop, supervised fine-tuning significantly improves performance across all other datasets and models, particularly for LLaMA 3.2-3B. Notably, LLaMA 3.2-3B even slightly outperforms o3-mini in political and religious datasets for attitude detection. Similarly, Qwen-14B surpasses o3-mini in political and religious datasets for value selection. These results demonstrate that SFT can effectively enhance the cultural intelligence of smaller models by distilling knowledge from larger models.
Qualitative analysis
We also conducted a qualitative analysis to examine the specific aspects in which SFT improves reasoning outputs. After manually reviewing 50 samples, we find out SFT mainly improve these three problems:
  • Inconsistency: Discrepancies between the reasoning process and the final answer.
  • Logical errors: Instances where the reasoning exhibits clear logical flaws, such as linking unrelated concepts.
  • Overlooking details: Cases where conclusions are drawn from general observations but important details in the conversation are missed.
We provide detailed examples for each category in Appendix B.2.
@@ 4 Related Work
Culture-aware LLMs
Culture-aware LLMs refer to models that account for cultural differences across various countries. Prior research has explored the cultural personas of LLMs and their consistency (Kharchenko et al., 2024; Rozen et al., 2024; Yao et al., 2024; Johnson et al., 2022; Saha et al., 2025). Zhong et al. (2024) show that the choice of prompting language can significantly influence the values expressed by LLMs. To enhance cultural awareness, researchers have proposed both single-culture and universally culture-aware LLMs by augmenting datasets and refining alignment strategies (Nguyen et al., 2023; Lin & Chen, 2023; Abbasi et al., 2023; Li et al., 2024a). To enhance cultural knowledge of LLMs, Myung et al. (2024) introduce a benchmark designed to help LLMs acquire cultural knowledge from everyday scenarios and Shi et al. (2024) collects unusual social norms from reddit. These culture-aware models have been shown to improve performance on culturally sensitive tasks, such as hate speech detection (Li et al., 2024a).
Value understanding
Value understanding is crucial in human-LLM interactions. Previous research has primarily focused on social norms and general values, with most studies addressing value detection in short contexts (Ren et al., 2024; Kiesel et al., 2023; Zhang et al., 2024). These works typically frame value detection as a classification task using entailment-like methods on short sentences, often without explicitly addressing cultural values. Fung et al. (2022) introduce a method for extracting social norms from conversations, emphasizing norm mining rather than cultural value identification. In contrast, our work focuses on understanding cultural values within long, real-life conversations, contributing to the development of culture-aware LLMs in the field of human-AI interaction.
@@ 5 Conclusion
We introduce CQ-Bench, a benchmark for evaluating LLMs’ ability to infer implicit cultural values in conversations. Unlike prior work that focuses on whether LLMs possess cultural knowledge or interpret value from a short text, CQ-Bench assesses their cultural reasoning through attitude detection, value selection, and value extraction. Our findings reveal that LLMs struggle with nuanced cultural intelligence. While larger models perform better, even state-of-the-art systems face challenges in cultural understanding. Notably, fine-tuning on just 500 examples significantly improves smaller models, demonstrating that cultural reasoning can be efficiently distilled with minimal data. CQ-Bench exposes gaps in LLMs’ cultural adaptability and serves as a foundation for advancing culturally intelligent AI. Future work can build on this to enhance LLM alignment with diverse human values.
6 Limitations
@@ A.1 Value set
We adopt seed values from the World value survey and Global Opinion. Each value has different option candidates, we provide possible option sets in Table 2
@@
Do you think that your country’s government should or should not have the right to do the following: Keep people under video surveillance in public areas? Definitely should not have the right Probably should not have the right Probably should have the right Definitely should have the right
In your view, how often do the following things occur in this country’s elections: Journalists provide fair coverage of elections? Very often Not often Not at all often
Work is a duty towards society. Agree Neither agree nor disagree Disagree
Apart from weddings and funerals, about how often do you pray? Frequently Occasionally Never
Having a strong leader who does not have to bother with parliament and elections. Very good Very bad
How important is it for people to help others? Important Not important
@@ A.2 Story generation
For each story, we will randomly predefined a scenario from those locations: company, school, neighborhood, national park, restaurant, amusement park, and airplane.
We remove very short stories (less than 400 words). The length of stories ranges from 500 to 900 words.
Story generation prompt is shown in Figure C.
@@ A.3 Story validation
We conduct 3 validations: incorporation check, consistency check and implicitness check. The prompts are shown in Figure C, C, C, C, C and C. For incorporation check and consistency check, which are directly related to the faithfulness of the dataset, we conduct human annotation on a small subset of 50 stories. We use Cohen’s Kappa McHugh (2012) to calculate inter-annotator agreement. The agreement score for the incorporation check is 0.904. Since we remove missing values from the ground truth, the remaining values are those the model considered as incorporated. Among these, human annotators agree with 98.2% of them.
The agreement score for the consistency check is lower, at 0.3, because the model applies a stricter criterion for consistency. For values marked as inconsistent, we exclude them from attitude detection, as they may confuse the model during attitude prediction. Similarly, during value selection, we do not include statements expressing an opposite attitude in the candidate options, as they could interfere with the model’s judgment. For values judged as consistent by the model, human annotators agree with 87.4% of them. We do not conduct consistency checks for the multiple attitude setting.
Value Original Speech Rewritten Speech
In your view, how often do the following things occur in this country’s elections: Voters are bribed?–Very often Speech: Voter bribing happens so often, it feels like a norm at this point. Explanation: The speech basically rephrase the value without providing supporting evidence But honestly, I wasn’t surprised. You hear about the rumors and tales around voting—sometimes it seems like it’s just part of the landscape now.
Table 3: An example of Implicitness check and rewriting.
Random Political Social Religious Ethical Multiple
AD 1665 301 213 425 270 1540
VS 2099 402 285 335 351 -
Table 4: Total datapoints for Attitude detection and total values for Value selection tasks.
@@
Appendix B Results
B.1 Category-specific results
The no reasoning results are shown in Table 3. The zero-shot and one-shot reasoning of smaller models are shown in 4. We show the rest of the results in Table 6. The social category dataset includes a single set of options: agree, disagree, and neither agree nor disagree. To study how well models understand the middle stance, we first remove questions where the ground truth is neither agree nor disagree, which eliminates about one-third of the data. We also remove the neither agree nor disagree option from the remaining examples, resulting in a fully binary dataset. On this binary version, o3-mini achieves an accuracy of 0.911. When we reintroduce the neither agree nor disagree option, performance drops to 0.811. However, when evaluated on the full dataset—including questions with neither as the correct answer—o3-mini only achieves 0.700 accuracy. This suggests that the inclusion of a middle stance can significantly challenge the model’s judgment.
@@ B.1 Category-specific results
Table 6: Results on category-specific dataset under zer-shot reasoning setting.
@@ B.2 Distillation on smaller models
We use reasoning results generated by o3-mini to fine-tune smaller models. We use the LLamafactory framework Zheng et al. (2024) and LoRA to accelerate the fine-tuning Hu et al. (2021). We train 5 epochs with a GPU of A6000 for 2-3 hours.
Qualitative analysis
@@
We find out the reasoning without fine-tuning has several issues. The first is inconsistency, for example, the reasoning mentioned the usage of alcohol multiple times, but the final answer fail to include values related to alcohol. Or the reasoning does not mention divorce at all, but the final answer choose some values related to divorce. The second is logical errors, for example “Oliver’s preference for a traditional family setup for raising kids matches When jobs are scarce, employers should give priority to people of this country over immigrants–Neither agree nor disagree ”. There is no direct logic between two statements. The third is overlooking details. For example, the one-shot reasoning is “The discussion also involves ethical considerations regarding the use of animals and the environment, which is reflected in the option :Hunting animals as a sport should be banned.–disagree. ” The reasoning itself sounds reasonable, however, the gold label is “Using animals for entertainment like in zoos/circuses is ethical.–disagree”. The SFT reasoning is “Finally, Harold’s statement on using animals for entertainment, which he finds shocking, aligns with the value that such practices are unethical.” The difference between two reasoning is that the latter one can catch the details of using animals for entertainment, while the former one only catch the use of animals and the environment.
Appendix C Prompts

### v2

@@ 4.3 CQ across Different LLMs
Although the AD task should be easier than value selection, its scores are even lower. One reason is that models struggle to distinguish neutral stances, such as neither agree nor disagree. While they can easily differentiate between agree and disagree, the presence of a neutral option can cause confusion. Even when a model correctly identifies agree, it may be distracted by the neutral choice and incorrectly select neither agree nor disagree. Additionally, models find it challenging to differentiate between varying levels of severity, such as not often and not at all often. Although it is also challenging for humans, humans can do better in identifying nuanced attitudes than models. We present results after merging options of varying levels of severity in Table 2.
Smaller models fail in long CoT reasoning in cultural intelligence task.
Although CoT reasoning significantly enhances performance in large models, smaller models often struggle with long CoT reasoning. LLaMA models, in particular, perform poorly in adhering to CoT reasoning across both AD and VS tasks. While Qwen 7B follows CoT reasoning well in AD, its performance declines significantly in VS as the reasoning steps become longer. A manual inspection reveals that its final output often consists of random values or irrelevant phrases that fail to focus on the given options. The DS-distill models exhibit slightly better instruction-following capabilities, outperforming Qwen 7B and LLaMA 8B in VS. While they do not strictly adhere to the prescribed format, their reasoning process generally aligns with the ideas provided in the prompt. However, their CQ reasoning ability remains weaker than their mathematical reasoning skills, resulting in final scores that are still lower than in the no-reasoning setting.
@@
VE tasks require strong reasoning and summarization capabilities, and smaller models often struggle with this task, frequently producing nonsensical outputs. As a result, we focus our evaluation on 5 models: Qwen 2.5 32B, GPT-4o-mini, o3-mini, o4-mini, and DeepSeek-V3. We also evaluate a smaller subset of 25 stories using o1, o3 and DeepSeek-R1. Interestingly, unlike attitude detection and value selection—where larger models consistently outperform smaller ones—we observe that a weaker model (Qwen 2.5 32B) can outperform stronger ones like o3-mini and o1. One possible explanation is that current CoT reasoning methods are not well-suited for open-ended generation; models tend to perform better when given predefined options.
4.4 CQ Across Different Value Categories
@@
We present results on category-specific datasets in Figure 3, showing the performance of six models in the no-reasoning setting. Complete results across all models and settings are provided in the Appendix B.1. Overall, the results show that models perform better in understanding political, social, and ethical values, achieving performance that is better than or comparable to the random setting. However, they perform worse in the religious values domain across both tasks. While models show stronger performance in attitude detection on the political dataset, they perform better in value selection on the ethical dataset. This suggests that it is easier for models to infer people’s political stances but more challenging for them to identify the specific topics being discussed in political contexts. In contrast, in the ethics domain, models find it easier to identify the main topics being discussed. For value extraction, overall, models perform better on category-specific results except for Deepseek-V3, as they only need to cover a single topic.
4.5 CQ ability learning for small models.
Attitude Detection Value Selection
Ethical Political Religious Social Ethical Political Religious Social
Model ZS GRPO SFT ZS GRPO SFT ZS GRPO SFT ZS GRPO SFT ZS GRPO SFT ZS GRPO SFT ZS GRPO SFT ZS GRPO SFT
Qwen 2.5 7B 0.652 0.696 0.704 0.708 0.731 0.718 0.645 0.515 0.713 0.657 0.662 0.714 0.357 0.754 0.690 0.399 0.639 0.572 0.249 0.515 0.519 0.395 0.662 0.517
Qwen 2.5 14B 0.644 0.696 0.707 0.701 0.731 0.731 0.649 0.704 0.708 0.699 0.668 0.718 0.674 0.800 0.687 0.574 0.707 0.681 0.440 0.527 0.552 0.567 0.702 0.608
Qwen 3 4B 0.548 0.641 0.607 0.724 0.728 0.714 0.545 0.631 0.588 0.633 0.624 0.624 0.498 0.709 0.745 0.555 0.678 0.667 0.281 0.512 0.556 0.468 0.635 0.659
Llama 3.2 3B 0.407 0.630 0.674 0.419 0.646 0.728 0.454 0.645 0.708 0.418 0.606 0.681 0.217 0.648 0.560 0.078 0.625 0.576 0.077 0.466 0.504 0.109 0.573 0.632
Llama 3.1 8B 0.596 0.700 0.704 0.638 0.748 0.721 0.574 0.730 0.678 0.574 0.709 0.695 0.304 0.751 0.645 0.261 0.695 0.576 0.164 0.642 0.504 0.280 0.708 0.632
O3-mini 0.728 - - 0.727 - - 0.703 - - 0.700 - - 0.837 - - 0.667 - - 0.531 - - 0.693 - -
Table 3: F1 scores for Attitude Detection and Value Selection across categories, models, and methods. ZS stands for zero-shot. We compare finetuned models with zero-shot prompting method.
In Section 4.3, we observe that smaller models struggle with long CoT reasoning, often failing to produce complete and coherent reasoning traces. We find that one-shot prompting does not notably improve the performance (Appendix B.2). To strengthen smaller models’ CQ reasoning ability, we explore two complementary fine-tuning approaches: Supervised Fine-Tuning (SFT) and Reinforcement Fine-Tuning (RFT). Both aim to improve the model’s ability to reason over cultural contexts but differ in learning strategy. SFT provides explicit reasoning supervision and encourages models to follow structured reasoning chains distilled from a larger model, while RFT promotes open-ended reasoning by allowing models to explore reasoning paths guided by reward signals.
For SFT, we distill reasoning traces from o3-mini using a random-setting dataset of 500 samples and fine-tune 5 smaller models (Qwen-2.5-7B, Qwen-2.5-14B, Qwen-3-4B, LLaMA 3.1-8B, and LLaMA 3.2-3B) with LoRA (Hu et al., 2021) for five epochs. Evaluation on category-specific datasets—which contain unseen values and thus approximate out-of-domain settings—reveals that SFT substantially improves reasoning quality and generalization across most domains. Notably, LLaMA 3.2-3B even surpasses o3-mini in political and religious AD tasks, while Qwen 14B exceeds it in the same domains for value selection (VS). We conduct a qualitative analysis to identify how SFT improves reasoning quality. Manual inspection of 50 samples reveals that SFT primarily mitigates issues of inconsistency between reasoning and answers, logical errors, and overlooked details. Details are provided in Appendix B.2.
For RFT, we adopt the GRPO method Shao et al. (2024), which achieves modest gains on the AD task and sometimes slightly lower than the SFT results, likely due to its simpler multiple-choice nature. However, it generally yields larger improvements on the more challenging VS task, which requires complex reasoning, as shown in Table 3. Implementation details are provided in Appendix B.2. Overall, both SFT and RFT significantly outperform zero-shot prompting, particularly on tasks involving long CoT reasoning.
@@ 5 Related Work
Culture-aware LLMs.
Culture-aware LLMs account for cross-cultural differences. Prior work has examined their cultural personas and consistency (Kharchenko et al., 2024; Rozen et al., 2024; Yao et al., 2024; Johnson et al., 2022; Saha et al., 2025b), showing that prompting language influences expressed values (Zhong et al., 2024). To improve cultural awareness, researchers have proposed both single- and multi-culture models through dataset augmentation and alignment (Nguyen et al., 2023; Lin and Chen, 2023; Abbasi et al., 2023; Li et al., 2024a; Wu et al., 2025). New benchmarks and datasets further support cultural knowledge acquisition (Myung et al., 2024; Shi et al., 2024), improving performance on tasks like hate speech detection (Li et al., 2024a).
Value understanding.
Value understanding is key to effective human-LLM interaction. Prior work has focused on detecting general social norms in short texts, often using classification or entailment-based methods (Ren et al., 2024; Kiesel et al., 2023; Zhang et al., 2024; Li et al., 2023). Wu et al. (2025) introduced an English dataset with 3,060 humancrafted multi-turn scenarios across 60 countries and found model challenges in interpreting nuanced cultural contexts. Others investigated datasets such as QA pairs and conversations and story scenarios in Persian Monazzah et al. (2025), Japanese & Sundanese Pranida et al. (2025), and Korean Kim et al. (2024), reporting model limitations in comprehending cultural and linguistic insights. Fung et al. (2022) also introduces a method for extracting social norms from conversations, emphasizing norm mining rather than cultural value identification. In contrast, our work focuses on understanding cultural values within long, real-life conversations, contributing to the development of culture-aware LLMs in human-AI interaction.
@@ 6 Conclusion
We introduce CQ-Bench, a benchmark for evaluating LLMs’ ability to infer implicit cultural values in conversations. CQ-Bench assesses model’s cultural intelligence through attitude detection, value selection, and value extraction. Our findings show that LLMs, including state-of-the-art models, struggle with nuanced cultural understanding. Fine-tuning on just 500 examples notably boosts smaller models, suggesting cultural reasoning can be efficiently distilled. CQ-Bench exposes gaps in LLMs’ cultural adaptability and serves as a foundation for advancing culturally intelligent AI. Future work can build on this to enhance LLM alignment with diverse human values.
Limitations
@@ A.1 Value Set
We adopt seed values from the World value survey(WVS) and Global Opinion. WVS is a global research project that explores people’s values and beliefs and how they change over time. The values cover different topics from personal beliefs to political stance. The results include over 200 values and 100 countries. Each value has different option candidates and we provide possible option sets in Table 4.
@@
Do you think that your country’s government should
or should not have the right to do the following:
Keep people under video surveillance in public areas?
Definitely should not have the right
Probably should not have the right
Probably should have the right
Definitely should have the right
In your view, how often do the following things occur
in this country’s elections: Journalists provide fair
coverage of elections?
Very often
Not often
Not at all often
Work is a duty towards society.
Agree
Neither agree nor disagree
Disagree
Apart from weddings and funerals, about how often
do you pray?
Frequently
Occasionally
Never
Having a strong leader who does not have to bother
with parliament and elections.
Very good
Very bad
How important is it for people to help others?
Important
Not important
@@ A.2 Story Generation
For each story, we will randomly predefined a scenario from those locations: company, school, neighborhood, national park, restaurant, amusement park, and airplane. We remove very short stories (less than 400 words). The length of stories ranges from 500 to 900 words.
@@ A.3 Story Validation
We conduct three validations: incorporation check, consistency check and implicitness check. We show the human annotation guidelines in Figure 4, 5, 6 and 7.
Figure 4: Screenshot of human annotation guideline (1)
Figure 5: Screenshot of human annotation guideline (2)
Figure 6: Screenshot of human annotation guideline (3)
Figure 7: Screenshot of human annotation guideline (4)
@@
Value Original Speech Rewritten Speech
In your view, how often
do the following things occur
in this country’s elections:
Voters are bribed?–Very often
Speech:
Voter bribing happens so often,
it feels like a norm at this point.
Explanation: The speech basically
rephrase the value without
providing supporting evidence
But honestly, I wasn’t surprised.
You hear about the rumors and
tales around voting—sometimes
it seems like it’s just
part of the landscape now.
Table 6: An example of implicitness check and rewriting.
@@
Table 7: Results on category-specific dataset under zero-shot reasoning setting.
Appendix B Results
B.1 Category-Specific Results
The no reasoning results are shown in Figure 3. The zero-shot and one-shot reasoning of smaller models are shown in Figure 8. We show the rest of the results in Table 7. The social category dataset includes a single set of options: agree, disagree, and neither agree nor disagree. To study how well models understand the middle stance, we first remove questions where the ground truth is neither agree nor disagree, which eliminates about one-third of the data. We also remove the neither agree nor disagree option from the remaining examples, resulting in a fully binary dataset. On this binary version, o3-mini achieves an accuracy of 0.911. When we reintroduce the neither agree nor disagree option, performance drops to 0.811. However, when evaluated on the full dataset—including questions with neither as the correct answer—o3-mini only achieves 0.700 accuracy. This suggests that the inclusion of a middle stance can significantly challenge the model’s judgment.
@@ B.2 Distillation on Smaller Models
We use reasoning results generated by o3-mini to fine-tune smaller models. We use the LLamafactory framework Zheng et al. (2024) and LoRA to accelerate the fine-tuning Hu et al. (2021). We train 5 epochs with a GPU of A6000 for 2-3 hours. The rank of LoRA is 8, and the learning rate is 0.0001. We try multiple combinations of hyperparameters before we choose the final hyperparameter set. All the experiments are single runs.
@@
We find out the reasoning without fine-tuning has several issues. The first is inconsistency, for example, the reasoning mentioned the usage of alcohol multiple times, but the final answer fail to include values related to alcohol. Another reasoning does not mention divorce at all, but the final answer choose some values related to divorce. The second is logical errors, for example: “Oliver’s preference for a traditional family setup for raising kids matches When jobs are scarce, employers should give priority to people of this country over immigrants.–Neither agree nor disagree.” There is no direct logic between two statements. The third is overlooking details. For example, the one-shot reasoning is “The discussion also involves ethical considerations regarding the use of animals and the environment, which is reflected in the option: Hunting animals as a sport should be banned. –Disagree.” The reasoning itself sounds reasonable, however, the gold label is “Using animals for entertainment like in zoos/circuses is ethical.–Disagree.” The SFT reasoning is “Finally, Harold’s statement on using animals for entertainment, which he finds shocking, aligns with the value that such practices are unethical. The difference between the two reasoningg is that the latter one identified details of using animals for entertainment, while the former one only identified the use of animals and the environment.
B.3 GRPO Fine-Tuning
We apply GRPO to fine-tune smaller LLMs on both attitude detection and value selection tasks. For the attitude detection task, which is formulated as multiple-choice classification, the reward signal is binary—1 for correct predictions and 0 otherwise. In contrast, the value selection task is a multi-label setting, where predictions can be partially correct. To account for this, we define the reward as the ratio of true positives to the total number of predicted values. If the number of selected values differs from the ground truth, the reward is set to 0.
Since partial rewards tend to slow convergence compared to binary signals, we train models until performance stabilizes. Empirically, we find that five epochs are sufficient for the attitude detection task, while value selection typically requires around fifteen epochs to converge. We set the rollout number to 16 to ensure stable and fair gradient updates from group-based advantage estimation. We train the model with 4 A100 for 2-3 hours.
B.4 Ablation Study
We conduct ablation studies using the Llama 3.2 3B model on both the Attitude Detection (AD) and Value Selection (VS) tasks. We fine-tune the model with {50, 100, 250, 500} stories, both with and without chain-of-thought (CoT) reasoning. Each configuration is trained with three random seeds, and we report the mean and standard deviation. To examine model size effects, we also evaluate the Qwen 2.5 14B model on the VS task.
Stories Setting Religious Social Ethical Political Multiple
50 W/O CoT 0.5474 (0.0061) 0.6369 (0.0058) 0.5025 (0.0126) 0.6965 (0.0095) 0.5835 (0.0043)
CoT 0.5537 (0.0156) 0.6761 (0.0101) 0.5074 (0.0212) 0.6910 (0.0141) 0.5729 (0.0072)
100 W/O CoT 0.6737 (0.0062) 0.6651 (0.0089) 0.6655 (0.0046) 0.6921 (0.0063) 0.5903 (0.0060)
CoT 0.6510 (0.0125) 0.6636 (0.0059) 0.6308 (0.0122) 0.6855 (0.0063) 0.5714 (0.0105)
250 W/O CoT 0.6792 (0.0099) 0.6917 (0.0059) 0.6951 (0.0182) 0.6955 (0.0042) 0.5617 (0.0117)
CoT 0.6463 (0.0111) 0.6479 (0.0102) 0.6383 (0.0257) 0.6689 (0.0110) 0.5615 (0.0052)
500 W/O CoT 0.6863 (0.0073) 0.7232 (0.0042) 0.6716 (0.0087) 0.6885 (0.0022) 0.5632 (0.0035)
CoT 0.6820 (0.0190) 0.6918 (0.0079) 0.6728 (0.0076) 0.7178 (0.0142) 0.5623 (0.0065)
Table 8: Ablation study on Attitude Detection using Llama 3.2 3B. We report average accuracy and standard deviation across three seeds.
Stories Setting Religious Social Ethical Political
50 W/O CoT 0.3204 (0.0070) 0.3673 (0.0177) 0.4355 (0.0132) 0.3571 (0.0339)
CoT 0.2849 (0.0035) 0.2921 (0.0145) 0.3326 (0.0074) 0.3116 (0.0116)
100 W/O CoT 0.3312 (0.0082) 0.4515 (0.0065) 0.5080 (0.0666) 0.4635 (0.0147)
CoT 0.3606 (0.0359) 0.4249 (0.0191) 0.5003 (0.0264) 0.3837 (0.0016)
250 W/O CoT 0.4668 (0.0263) 0.5429 (0.0119) 0.6525 (0.0100) 0.5666 (0.0057)
CoT 0.4172 (0.0128) 0.5300 (0.0051) 0.6304 (0.0132) 0.5263 (0.0271)
500 W/O CoT 0.4881 (0.0113) 0.5859 (0.0134) 0.7076 (0.0147) 0.5957 (0.0110)
CoT 0.4036 (0.0059) 0.5672 (0.0170) 0.6089 (0.0100) 0.5181 (0.0029)
Table 9: Ablation study on Value Selection using Llama 3.2 3B. We report average accuracy and standard deviation across three seeds.
Stories Setting Religious Social Ethical Political
50 W/O CoT 0.4272 (0.0156) 0.6093 (0.0087) 0.6916 (0.0079) 0.5625 (0.0154)
CoT 0.4642 (0.0091) 0.5971 (0.0179) 0.6758 (0.0093) 0.5996 (0.0021)
100 W/O CoT 0.4210 (0.0167) 0.5909 (0.0067) 0.6902 (0.0179) 0.5708 (0.0062)
CoT 0.4951 (0.0109) 0.5876 (0.0101) 0.6720 (0.0137) 0.5634 (0.0251)
250 W/O CoT 0.4099 (0.0157) 0.6002 (0.0060) 0.6794 (0.0117) 0.5821 (0.0149)
CoT 0.5461 (0.0076) 0.5868 (0.0148) 0.6697 (0.0148) 0.5962 (0.0082)
500 W/O CoT 0.4817 (0.0069) 0.6129 (0.0067) 0.7672 (0.0043) 0.6663 (0.0032)
CoT 0.5465 (0.0122) 0.6991 (0.0148) 0.7524 (0.0200) 0.6935 (0.0039)
Table 10: Value Selection results on Qwen 2.5 14B.
We summarize key findings as follows:
  • Impact of dataset size. In attitude detection, increasing dataset size does not strongly correlate with performance, likely because the task involves only two to three attitude options and the pattern is quickly learnable. In contrast, value selection shows consistent improvement as dataset size increases, especially from 50 to 250 samples.
  • Effect of CoT reasoning. Surprisingly, non-reasoning (W/O CoT) fine-tuning often outperforms CoT in the Llama 3.2 3B model, particularly in the VS task. Manual inspection reveals that smaller models struggle to maintain coherent reasoning chains, often producing disorganized outputs.
  • Effect of model size. For the Qwen 14B model, CoT begins to show advantage as dataset size increases, indicating that larger models can better leverage reasoning supervision. While the improvements are not universally significant, results suggest that CoT becomes beneficial only beyond a certain data scale and model capacity.
  • Alignment with prior findings. Our observations are consistent with prior work (kojima2023cotnotcot), which shows that CoT mainly benefits symbolic or mathematical reasoning tasks, and may not generalize as effectively to social or value-based reasoning.
B.5 Case study on value expression
To better assess the model’s ability to express values implicitly, we conduct a small case study using Qwen 2.5 7B. The model is prompted to respond to an ongoing conversation given a target value, with the goal of reflecting the value implicitly while maintaining natural conversational flow. We compare the vanilla Qwen 2.5 7B model with its GRPO-tuned counterpart.
To evaluate the quality of generated speech, we employ an LLM-as-a-judge framework that assesses three dimensions: value reflection, speech naturalness, and value implicitness. We experiment with two judging strategies: (1) pairwise comparison, where the LLM directly compares two speeches, and (2) independent rating, where it scores each speech separately before comparison. Across both evaluation methods, the GRPO-tuned model performs slightly better, with approximately 5% of speeches showing improvement, though the gain is not statistically significant. This case study suggests that while GRPO tuning enhances implicit value expression modestly, further improvements could be achieved by jointly strengthening detection and expression capabilities.
Appendix C Human Annotations
To compare the human performance with LLMs, we ask several human participants from CS undergraduates and master’s students. We cover students from different background including 2 Chinese students and 2 Indian students. We provide the exact prompt we use for LLM to those students.
Appendix D Prompts